OpenStudy (anonymous):

There is a value of a such that subtracting a-4 from 4a+16 gives an answer of -25. What is that value of a? (We suggest that you check your answer.)

OpenStudy (anonymous):

4a+16 -(a-4) = -25

OpenStudy (anonymous):

\((4a + 16) - (a-4) = -25\) \(3a + 20 = -25\) Find a..

OpenStudy (anonymous):

Firstly subtract 20 from both the sides and then divide by 3 both the sides.. :)

OpenStudy (anonymous):

ok